Compute algebraically the resultant of the

following coplanar forces: 100 N at 30O

, 141.4


N at 45O

, and 100 N at 240O


. Check your


answer graphically.


1
Expert's answer
2022-03-20T18:51:05-0400

"F_x=F_1\cos\alpha_1+F_2\cos\alpha_2+F_3\cos\alpha_3=136.6~N,"

"F_y=F_1\sin\alpha_1+F_2\sin\alpha_2+F_3\sin\alpha_3=63 .4~N,"

"F=\sqrt{F_x^2+F_y^2}=150.6~N,"

"\phi=\arctan\frac{F_y}{F_x}=25°."